OpenStudy (anonymous):

Evaluate the formula. A=4pir^2 if pi=3.14 and r =18 in. The value of A is?

OpenStudy (legends):

4*3.14*18*2? like that

OpenStudy (legends):

@Jamie1019

OpenStudy (anonymous):

Yes, we multiply 4 * 3.14 * 18^2

OpenStudy (anonymous):

Actually that's 18^2 @Legends which is the same thing as 18 * 18

OpenStudy (legends):

got it ok

OpenStudy (anonymous):

So we have: \(4 \times 3.14 \times 18 \times 18\) Plug that in your calculator @Jamie1019

OpenStudy (anonymous):

4069.44

OpenStudy (anonymous):

Yes, that is our answer.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Response'.

sammixboo (sammixboo):

^ http://goo.gl/zB9APU